CRY+VIS Interface Meeting
Date: 2022/5/13 13:00-13:30
Participants: Takahashi, Sato, Ushiba, Aso, Ikeda, Akutsu, Miyoki, Hirata, Node, Kimura, Yano
- EX: Preparation for the cooling of EXC.
- IY: Finalization. Health check. ADC noise. Go to implementation of SI.
- EY: Finalization. Health check. ADC noise. BF LVDT noise. Cabling repair (P7-2). Go to implementation of SI.
- Type-B: Calibration of geophone in SR2. Stand-alone SM driver for SR3.
- Type-Bp: FC onto PRM (small regidual FC on AR). Traverser limit SW: Installation in PRM in the early next week. MSB: Installation in PRM in the late next week. FC onto PR3 on 5/23.
- Type-C: OMC installation from 5/23. Health check of OMMT1/2/OSTM.
- Replacement of flip cables: Replaced for TMS in EY.
- TM height setting:
- Oplev: Windshield in the next.
- Commissioning in air:
- Commissioning in vacuum:
- Acceptance check: Review from 5/13.
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: How to treat thermal load?
- Heater control: Pending.
- TMS: Thermometer
Cryo-shield: Finnese measurement, height check --> Cooler ON.
Date: 2022/5/13 14:00-14:40
Participants: Takahashi, Ushiba, Fabian, Aso, Tanaka, Washimi
- Replaced the main card of the LVDT driver for the BF damper in EY.
- Repaired the cabling for the thermometer in EY.
- Closed four feedthrough flanges in EY and two ones in IY.
- Replaced the HPCD for the IP in IY
I kept working on the geophones calibration transfer functions. See slides.
We applied FC to PRM.
- Measured the transfer functions of suspensions for OMMT1, OMMT2, and OSTM. Adjusted the position of H2 OSEMs for OMMT2/OSTM, and H4 OSEM for OMMT2. The baffle plates on the suspension frames were removed temporally, and they were returned back.
- Type-A: Review of IY and EY.
- Type-B/Bp/C: Installation of traverse limit switches and MSB in PRM. Calibration of the geophones in SR2.
(Date in Kamioka)
The week from 5/16
- Takahashi: 5/17-20
- Sato: 5/16-20
- Hirata: 5/17-20
The week from 5/9
- Takahashi: 5/9-13
- Sato: 5/9-13
- Hirata: 5/9-13